OpenStudy (anonymous):

What is the inverse of f(x) = 5x+4?

hero (hero):

f(x) = 5x+4 y = f(x), therefore, y = 5x + 4 y - 4 = 5x (y-4)/5 = x (x-4)/5 = y (x-4)/5 = f^(-1)(x)

OpenStudy (anonymous):

Original \[y = 5x + 4\] Solve for x \[x = (y-4)/5\] Making your inverse \[f^-1(x) = (x-4)/5\]

OpenStudy (anonymous):

Ah sorry Hero, Didn't see your response there.